OpenStudy (saifoo.khan):

Quick Question, Area under speed-time graph is the distance travelled?? T or F?

OpenStudy (anonymous):

t

OpenStudy (dumbcow):

True speed = distance/time *time = distance

OpenStudy (anonymous):

umm is this a physics question?

OpenStudy (saifoo.khan):

No, this is Math,

OpenStudy (anonymous):

okies :)

OpenStudy (saifoo.khan):

Now it's fine. :D

OpenStudy (anonymous):

Distance to me is an absolute value. No directional sense. So, as I pointed in the link and as MrYantho suggested, I like displacement.

OpenStudy (saifoo.khan):

Okay!! (: can u solve the question in the picture?

OpenStudy (anonymous):

Saifoo - sorry man....I did this stuff 22 years ago and it is hard to remember. I can google it up and get it. But, you or someone current is going to be a lot better at it.

OpenStudy (saifoo.khan):

No worries. Thanks for trying! :)

OpenStudy (anonymous):

sorries me cant do it either :(

OpenStudy (saifoo.khan):

np

OpenStudy (anonymous):

um its just at constant speed not instantaneous right?

OpenStudy (saifoo.khan):

first it has increasing speed, then constant speed, i want to draw the distance-time graph for this.

OpenStudy (anonymous):

could you wait a bit ill try to help this stuff is in my physics book

OpenStudy (anonymous):

its the math it uses

OpenStudy (saifoo.khan):

This might help,

OpenStudy (anonymous):

the formula you just use is a regular d=r/t

OpenStudy (saifoo.khan):

idk. this is the screenshot form the marking scheme.. i have no idea how they did that.

OpenStudy (anonymous):

formula for aceleration final speed= initial speed + speed change

OpenStudy (saifoo.khan):

\[a = \frac{v-u}{t}\]Right?

OpenStudy (saifoo.khan):

But how we will draw the speed-time? :l

OpenStudy (anonymous):

on you graph y is the speed x is the time

OpenStudy (saifoo.khan):

wait. sorry.. that is already drawn in the question.. i want displacement time. hehe,

OpenStudy (anonymous):

displacement ? umm i dont understand displacement from what?

OpenStudy (saifoo.khan):

read the question again, that might help you the part (i) http://assets.openstudy.com/updates/attachments/4f1a4dd4e4b04992dd22a68d-saifoo.khan-1327124195861-dscn4553.jpg

OpenStudy (anonymous):

it is cut off

OpenStudy (anonymous):

can you type the rest it?

OpenStudy (saifoo.khan):

The first part says to sketch the region whose area represents and stuff.. t And the cut off part says "find the value of T"

OpenStudy (anonymous):

first time change of 1 cyclist s=(v) (t) + 1/2 ( a)( t) use formula again to find displacement again of the cyclist

OpenStudy (anonymous):

then you may probably use d=rt to figure out the rest of the stuff i think this is how it is done :\

OpenStudy (anonymous):

it's true that the area under speed-time graph is distance you can get this from this formula d=vt